OpenStudy (anonymous):

CRAZY CALCULUS PROBLEM: a searchlight is located at point A, 40 ft from a wall. The search light revolves counterclockwise at a rate of (pi/30) radians per second. At any point B on the wall, the strength of the light L, is inversely proportional to the square of the distance d from A; that is, at any point on the wall L=(k/d^2). At the Closest point P, L=10,000 lumens.

OpenStudy (anonymous):

I will now draw a picture!

OpenStudy (anonymous):



OpenStudy (anonymous):

a.) Find the constant of proportionality k. <--- does that mean just solve for k????

hartnn (hartnn):

you have L = 10000 for d=40 just plug it in L=(k/d^2)

OpenStudy (anonymous):

Yep, just solve for \(k\).

hartnn (hartnn):

i am sure there are more parts of this problem yet to come... (because this isn't calculus yet :P)

OpenStudy (anonymous):

THat makes sense. Oh, the calculus will come soon enough. x) b.) express L as a function of theta, the angle formed by the lines AP and AB...

hartnn (hartnn):

can you find 'd' in terms of theta first ?? using trigonometry ? hint : cos ratio

hartnn (hartnn):

lot easier to find cos theta from that right triangle... cos theta =... ?

hartnn (hartnn):

or sec theta will also do

OpenStudy (anonymous):

sOrry,my computer froze. hold on, i'm trying to process xD

hartnn (hartnn):

i am sure you know how to find cos theta, right ? also no calculus yet :P i am curious about it being crazy!

OpenStudy (anonymous):

Gah! Leave me alone! haa jk DOn't go! Cos(theta)=40/d!

hartnn (hartnn):

so, d= 40/cos theta or d= 40 sec theta just plug this in L = k/d^2 and done with b)!

hartnn (hartnn):

not to mention that you already know 'k' from part a)

hartnn (hartnn):

i won't leave until i make you solve this.... and i am not jk :P

OpenStudy (anonymous):

lol thanks! So... does it reduce to 10,000cos(theta)?

hartnn (hartnn):

can you please post all the parts ? so that we can think on further steps.... first what you got 'k' from part a) ?

hartnn (hartnn):

did you forget to square cos theta ? because i get L = 10000 cos^2 theta

OpenStudy (anonymous):

Do you mean all the parts to the problem? c.) How fast (in Lumens/Second) is the strength of the light changing when theta=(pi)/4? Justify ura nswere! (ANd yeah, i forgot to square -_-)

hartnn (hartnn):

now comes the calculus. can you differentiate, L = 10000 cos^2 theta ?

OpenStudy (anonymous):

(with respect to t)

OpenStudy (anonymous):

shouldnt it be 400000Cos^2(theta). did u forget to square D in part a? cuz i did lol

hartnn (hartnn):

:O wait, let me recalculate

OpenStudy (anonymous):

^_^

hartnn (hartnn):

i still get L=10000 cos^2 theta k=16*10^6

hartnn (hartnn):

d^2 =40^2 = 1600

OpenStudy (anonymous):

yeah, so that is d^2, then it's multiplied by L. or am I doing this wrong?

hartnn (hartnn):

1000 = k/ 1600 k = 16*10^6 L = 16*10^6 / d^2 d= 40 sec theta L = 16*10^6 / (40^2 sec^2 theta ) L = (16*10^6 / 1600 ) cos^2 theta L =10000 cos^2 theta.

OpenStudy (anonymous):

Oops ^_^ lol! Sorry! I promise im not always this dumb! I'm one of the only 3 A's in my calculus class!!!! lol ok, go on xD

hartnn (hartnn):

any doubts so far ? can you differentiate, L = 10000 cos^2 theta with respect to theta ? its ok :)

OpenStudy (anonymous):

You want the derivative with respect to time.

hartnn (hartnn):

oh yes :P

OpenStudy (anonymous):

20000Cos(theta)*(-sin(theta))

OpenStudy (anonymous):

But you need it with respect to theta for the chain rule\[ \frac{dL}{dt}=\frac{dL}{d\theta}\frac{d\theta}{dt} \]

hartnn (hartnn):

that is dL/ d theta = 20000Cos(theta)*(-sin(theta)) can you simplify a bit ? and yes, you need dL/dt, so i suggest you divide both sides by dt

OpenStudy (anonymous):

"The search light revolves counterclockwise at a rate of (pi/30) radians per second. " This is a hint about \(d\theta/dt\)

OpenStudy (anonymous):

Simplify? Uhm... i dont know.

hartnn (hartnn):

\(2 \sin \theta \cos \theta = \sin 2\theta \) but not necessary.

hartnn (hartnn):

so, we have \(dL = -10000 \sin 2\theta \: d\theta \\dL/dt = -10000 \sin 2\theta \: (d\theta/dt) \) dL/ dt is what u need to find, and theta, dtheta/dt is given.

OpenStudy (anonymous):

ok. So how do I find dL/dt?

hartnn (hartnn):

theta, dtheta/dt is given.can you figure out those ?

OpenStudy (anonymous):

no :(

hartnn (hartnn):

what does this give you ? The search light revolves counterclockwise at a rate of (pi/30) radians per second''

hartnn (hartnn):

and part D directly gives you '..... when theta=(pi)/4'

OpenStudy (anonymous):

I'm not sure what I'm suppoased to be doing 0.0

hartnn (hartnn):

The search light revolves counterclockwise at a rate of (pi/30) radians per second----------> dθ/dt = pi/30, θ = pi/4, given in part d) of question, plug those in dL/dt=−10000sin2θ(dθ/dt) and you get dL/dt

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHH! BLAH! Lol ok! Hold on xD

OpenStudy (anonymous):

-1047.2?

hartnn (hartnn):

yes, that correct....or you can keep that as -1000pi/3

OpenStudy (anonymous):

Ok, so the strength of the light is increasing at pi/4? and why is it negative?

hartnn (hartnn):

when theta = pi/4 strength of light is changing at the rate of -1000pi/3 lumens/sec means decreasing @ 1000pi/3 lumens/sec (negative implies decreasing, positive implies increasing) means in one second, strength of light decreases by 1000pi/3 lumens

OpenStudy (anonymous):

I ment increasing xD lol Idk why i wrote decreasing thank you so much! i have no idea what time it is where u are but. GOOD NIGHT xD

hartnn (hartnn):

good night and sweet dreams to you ^_^ (its 1.07 pm here :) )

hartnn (hartnn):

welcome so much ! ^_^

OpenStudy (anonymous):

wahhh! 11:39. Goku has saved the day again